Erin's age is 3 times Warren's. In 4 years she will be twice as old as he will be. How old is each now? How would you check your answer?
Answer by checkley77(12844)

E=3X
(3X+4)=2(X+4)
3X+4=2X+8
3X-2X=8-4
X=4 AGE OF WARREN NOW.
E=3*4=12 AGE OF ERIN NOW.
PROOF:
(12+4)=2(4+4)
16=2*8
16=16